OpenStudy (anonymous):

please help me!!!! 5 2 5 -------- - ----- = ----- x^2-7x+12 3-x x-4

OpenStudy (anonymous):

please can someone help!!!!

OpenStudy (mertsj):

Factor the first denominator.

OpenStudy (anonymous):

(x-4)(x-3)

OpenStudy (mertsj):

For the second fraction, please note that: \[- \frac{2}{3-x}=+\frac{2}{x-3}\]

OpenStudy (mertsj):

Now, write each fraction with the common denominator (x-4)(x-3)

OpenStudy (anonymous):

how if the end of the equations doesnt have the same lcd

OpenStudy (anonymous):

\[\frac{5}{x^2-7 x+12}-\frac{2}{3-x}-\frac{5}{x-4}\neq 0 \]\[\frac{1}{x-3}\neq 0 \]

OpenStudy (anonymous):

so the answer is no solution?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ohhhh